# Statistical Guarantees of Correctness Coverage for Medical Multiple-Choice Question Answering

### v1

@@ 4.3 Sensitivity Analysis
The minimal performance discrepancy between the two methods highlights the robustness and practicality of employing the frequency method as a substitute for logit-based evaluations. This is particularly advantageous in applications where access to logit data is restricted, enabling a reliable and effective alternative for model performance assessment under such constraints.
Table 3: Results of the Auroc.
Datasets LLMs Frequency Logit MedMCQA Llama-3.2-1B-Instruct 0.5872 0.6197 Llama-3.2-3B-Instruct 0.6982 0.7005 Qwen2.5-1.5B-Instruct 0.6337 0.6517 Qwen2.5-3B-Instruct 0.6418 0.6930 MedQA Llama-3.2-1B-Instruct 0.5819 0.6041 Llama-3.2-3B-Instruct 0.7375 0.7484 Qwen2.5-1.5B-Instruct 0.5752 0.6241 Qwen2.5-3B-Instruct 0.5693 0.6609 MMLU Llama-3.2-1B-Instruct 0.6380 0.6358 Llama-3.2-3B-Instruct 0.7501 0.7657 Qwen2.5-1.5B-Instruct 0.7159 0.7771 Qwen2.5-3B-Instruct 0.6207 0.7784
Reliability Measurement
To evaluate model reliability, we conducted a comparative analysis using two distinct measurement strategies: frequency-based metrics and logit-based metrics, specifically focusing on the MedMCQA dataset. The assessment centered on two critical performance indicators: Empirical Error Rate (EMR) and Average Prediction Set Size (APSS).
Figure 5 presents the performance using frequency-based measurements, while Figure 6 illustrates the results with logit-based metrics. The logit-based approach demonstrates a consistent and gradual increase in both EMR and APSS across varying values of the risk level $\alpha$. Models evaluated with logit-based metrics exhibit more stable performance, with error rates and prediction set sizes increasing smoothly and predictably as $\alpha$ rises.
Conversely, the frequency-based approach, as shown in Figure 5, results in more pronounced fluctuations, particularly in EMR. The frequency-based method shows a steeper decline in average set size; however, the associated increase in EMR is more erratic when compared to the logit-based approach. This indicates that frequency-based metrics may lead to greater variability in model performance as the confidence threshold changes.
@@
Nevertheless, in practical scenarios where access to the model’s internal logits may not be feasible, frequency-based metrics offer a practical alternative. While these metrics may not achieve the same level of precision and stability as the logit-based approach, they still provide a reasonable approximation of performance and reliability. This adaptability is particularly advantageous in real-world applications where the internal model structure or logit data is not accessible, enabling robust reliability assessments using only observable model outputs.
Figure 6: Reliability measurement using logit-based metrics on the MedMCQA dataset, demonstrating stable performance with a smooth increase in EMR and APSS as $\alpha$ increases.
5 Conclusion
In this study, we present a novel risk control framework for medical MCQA tasks by integrating CP with self-consistency theory. Our approach not only preserves the model-agnostic and distribution-free advantages of CP but also extends its functionality beyond simple coverage control by introducing a monotonically decreasing loss function that effectively manages task-specific metrics. Extensive empirical evaluations conducted on three medical MCQA datasets, utilizing four LLMs, demonstrated that our framework consistently achieved user-specified miscoverage rates. Additionally, our analyses indicated a clear inverse relationship between the APSS and the associated risk level, highlighting the potential of APSS as a robust metric for quantifying LLM uncertainty. The proposed framework offers a statistically rigorous and interpretable method to enhance the trustworthiness of LLMs in high-stakes medical applications by mitigating risks linked to model hallucinations. Future research will investigate the adaptation of this framework to other domain-specific QA tasks and its integration with advanced LLM architectures to further improve reliability and transparency.

### v2

@@ 4.3.2 AUROC Analysis
The minimal performance discrepancy between the two methods highlights the robustness and practicality of employing the frequency method as a substitute for logit-based evaluations. This is particularly advantageous in applications with restricted access to logit data, enabling a reliable and effective alternative for model performance assessment under such constraints.
Table 3: Results of the AUROC metric.
Datasets LLMs Frequency Logit
MedMCQA Llama-3.2-1B-Instruct 0.5872 0.6197
Llama-3.2-3B-Instruct 0.6982 0.7005
Qwen2.5-1.5B-Instruct 0.6337 0.6517
Qwen2.5-3B-Instruct 0.6418 0.6930
MedQA Llama-3.2-1B-Instruct 0.5819 0.6041
Llama-3.2-3B-Instruct 0.7375 0.7484
Qwen2.5-1.5B-Instruct 0.5752 0.6241
Qwen2.5-3B-Instruct 0.5693 0.6609
MMLU Llama-3.2-1B-Instruct 0.6380 0.6358
Llama-3.2-3B-Instruct 0.7501 0.7657
Qwen2.5-1.5B-Instruct 0.7159 0.7771
Qwen2.5-3B-Instruct 0.6207 0.7784
4.3.3 Reliability Measurement
A comparative analysis was conducted to evaluate model reliability using two distinct measurement strategies: frequency-based metrics and logit-based metrics, explicitly focusing on the MedMCQA dataset. The assessment centered on two critical performance indicators: Empirical Error Rate (EMR) and Average Prediction Set Size (APSS).
Figure 5 presents the performance using frequency-based measurements, while Figure 6 illustrates the results with logit-based metrics. The logit-based approach demonstrates a consistent and gradual increase in EMR and APSS at varying risk levels $\alpha$. values. The proposed models evaluated with logit-based metrics exhibit more stable performance, with error rates and prediction set sizes increasing smoothly and predictably as $\alpha$ rises.
Figure 6: Reliability measurement using logit-based metrics on the MedMCQA dataset, demonstrates stable performance with a smooth increase in EMR and APSS as $\alpha$ increases.
Conversely, the frequency-based approach, as shown in Figure 5, results in more pronounced fluctuations, particularly in EMR. The frequency-based method shows a steeper decline in average set size; however, the associated increase in EMR is more erratic when compared to the logit-based approach. It is indicated that frequency-based metrics may lead to greater variability in model performance as the confidence threshold changes.
@@ 4.3.3 Reliability Measurement
Nevertheless, frequency-based metrics offer a practical alternative in practical scenarios where access to the model’s internal logits may not be feasible. While these metrics may not achieve the same level of precision and stability as the logit-based approach, they still provide a reasonable approximation of performance and reliability. This adaptability is advantageous in real-world applications where the internal model structure or logit data is inaccessible, enabling robust reliability assessments using only observable model outputs.
5 Conclusions
This study presents an enhanced CP framework to achieve risk control in medical MCQA tasks. By integrating self-consistency theory, the enhanced CP framework provides a statistically rigorous and interpretable method to improve the trustworthiness of LLMs in high-stakes medical applications by mitigating risks linked to model hallucinations. Comparative experiments demonstrated that the enhanced CP framework not only preserves CP’s model-agnostic and distribution-free advantages extends its functionality beyond simple coverage control by introducing a monotonically decreasing loss function that effectively manages task-specific metrics. Additionally, the enhanced CP framework consistently achieved user-specified miscoverage rates. The results analysis also indicated a clear inverse relationship between the APSS and the associated risk level, highlighting the potential of APSS as a robust metric for quantifying LLM uncertainty. Future research will investigate the adaptation of this framework to other domain-specific QA tasks and its integration with advanced LLM architectures to improve reliability and transparency further.